# Sliding Puzzles Gym: A Scalable Benchmark for State Representation in Visual Reinforcement Learning

### v1

@@ Abstract
Learning effective visual representations is crucial in open-world environments where agents encounter diverse and unstructured observations. This ability enables agents to extract meaningful information from raw sensory inputs, like pixels, which is essential for generalization across different tasks. However, evaluating representation learning separately from policy learning remains a challenge in most reinforcement learning (RL) benchmarks. To address this, we introduce the Sliding Puzzles Gym (SPGym), a benchmark that extends the classic 15-tile puzzle with variable grid sizes and observation spaces, including large real-world image datasets. SPGym allows scaling the representation learning challenge while keeping the latent environment dynamics and algorithmic problem fixed, providing a targeted assessment of agents’ ability to form compositional and generalizable state representations. Experiments with both model-free and model-based RL algorithms, with and without explicit representation learning components, show that as the representation challenge scales, SPGym effectively distinguishes agents based on their capabilities. Moreover, SPGym reaches difficulty levels where no tested algorithm consistently excels, highlighting key challenges and opportunities for advancing representation learning for decision-making research.
@@ 1 Introduction
Learning meaningful representations from raw sensory inputs, such as visual data, is fundamental to reinforcement learning (RL) agents’ ability to generalize across different tasks in complex, open-world environments. In visual RL, agents must process high-dimensional pixel data, extract useful features, and utilize these features for decision-making. This becomes especially crucial as real-world applications demand adaptability to unstructured and diverse observations. However, disentangling representation learning from other learning tasks, such as policy optimization or dynamics modeling, remains a key challenge in RL benchmarks.
Most RL benchmarks, including widely adopted ones like Atari [2], DeepMind Control Suite [16], and DeepMind Lab [1], entangle representation learning with other learning challenges, making it difficult to isolate the agent’s ability to form robust representations. More specialized benchmarks, such as the Distracting Control Suite [15], emphasize robustness to visual noise but focuses on filtering out irrelevant complexity.
We address this gap by introducing the Sliding Puzzles Gym (SPGym), [^1] Code available at https://github.com/bryanoliveira/sliding-puzzles-gym. a scalable RL benchmark designed to test agents’ visual representation learning capabilities. SPGym extends the 15-tile puzzle with variable grid sizes and observation spaces, such as image-based tiles, allowing for a controlled and scalable increase in representation complexity. Crucially, SPGym keeps the underlying environment dynamics and task structure constant, isolating the representation learning challenge.
We evaluate the performance of both model-free and model-based RL agents on SPGym, comparing their ability to generalize across increasingly complex visual input spaces. Our experiments reveal that, as representation learning complexity increases, SPGym effectively differentiates between agents based on their ability to form compositional and generalizable state representations. Specifically, model-based agents such as DreamerV3 [4] outperform model-free ones like PPO [14] in most scenarios, though even these methods struggle with higher levels of visual complexity.
Contributions
@@
  • We introduce the Sliding Puzzles Gym (SPGym), a scalable benchmark for assessing representation learning in visual decision-making. SPGym achieves this by gradually increasing the complexity of visual observations without changing the underlying task structure.
  • We provide an experimental evaluation of model-free and model-based RL agents, demonstrating the benchmark’s effectiveness in distinguishing between agents’ representation learning capabilities.
  • We reveal limitations in current RL methods’ ability to handle complex and diverse visual inputs, showcasing the scalability and challenge SPGym offers as a benchmark.
Our experiments show that SPGym offers a challenging testbed for visual representation learning in RL, making it a valuable tool for future research aimed at improving agents’ visual representation capabilities in open-world environments.
@@ 2 Related Work
Reinforcement learning benchmarks are essential tools for evaluating agents’ capabilities across diverse tasks. Popular benchmarks like Atari [2], DeepMind Control Suite [16], and DeepMind Lab [1] focus primarily on assessing policy learning and generalization. Yet, these benchmarks inherently combine representation learning with policy optimization and dynamics modeling, making it difficult to isolate the specific impact of representation learning on an agent’s overall performance. Even methods aimed at improving representation learning for RL agents, such as DARLA [5], CURL [8], RAD [9], DrQ [7], CBM [10], and CycAug [11] are typically evaluated within these entangled settings. While these methods improve agents’ ability to learn from raw visual inputs, they do not provide a clear, isolated assessment of their representation learning capabilities.
ProcGen [3] takes a step toward evaluating generalization through diversity by offering procedurally generated levels, challenging agents to adapt to unseen environments. However, like other benchmarks, it does not fully separate representation learning from policy and dynamics learning. Agents must simultaneously learn representations, policies, and environment dynamics, making it difficult to attribute performance improvements solely to better representations. Alternatively, the Distracting Control Suite [15], an extension of DM Control, introduces visual distractions to evaluate agents’ robustness to irrelevant variations. While it allows for parametrizable control over distraction complexity, its primary focus is on agents’ ability to ignore these distractions rather than to extract meaningful information from visually complex observations.
SPGym’s design sets it apart from existing benchmarks by providing a controlled environment where difficulty arises solely from the complexity of visual observations, rather than changes in task dynamics or objectives. This makes SPGym particularly well-suited for comparing RL algorithms that specifically target improvements in representation learning.
@@ 3 The Sliding Puzzles Gym
Figure 1: Overview of the Sliding Puzzles Gym (SPGym). The framework extends the 15-tile puzzle by incorporating image-based tiles, allowing scalable representation complexity while maintaining fixed environment dynamics.
The Sliding Puzzles Gym (SPGym) extends the 15-tile puzzle, where players rearrange a shuffled $4\times 4$ grid of numbered tiles into the correct sequence using four actions: UP, DOWN, LEFT, or RIGHT. SPGym generalizes this concept by supporting variable grid sizes, ranging from $2\times 2$ to larger $H\times W$ configurations, and incorporating diverse observation spaces. In this work, we focus on $3\times 3$ grids to emphasize the challenge of visual representation learning using image-based observations.
SPGym’s internal representation is a 2D array of tile indices representing the current grid state. We provide wrappers that convert these states into agent-friendly observations, such as one-hot encodings or, as used in this paper, image overlays. Figure 1 presents the main components of SPGym.
At each time step, the reward is computed as
$r_{t}=\begin{cases}-\frac{\sum_{i,j}|x_{i,j}-x^{*}_{i,j}|+|y_{i,j}-y^{*}_{i,j}|}{D},&\text{if action is valid}\\ -1,&\text{if action is invalid}\\ +10,&\text{if puzzle is solved}\end{cases}\text{,}$ (1)
with
@@
SPGym provides two methods for generating the puzzle’s initial state. The first method, used in this paper, generates a random $H\times W$ array and adjusts the parity of the puzzle by swapping the first two tiles if necessary to ensure solvability [6]. The second method, which can support curriculum learning in future work, starts with a solved puzzle and applies a random series of valid moves to generate an initial state, enabling easier starting puzzle configurations.
Wrappers and Observation Modalities
SPGym offers flexible wrappers that transform internal states into different modalities, ranging from simple one-hot encodings to complex visual observations, such as real-world images and text (see Figure 2). This flexibility makes SPGym adaptable for exploring how reinforcement learning agents handle diverse observation spaces, enhancing its utility as a benchmark for representation learning. In this paper, however, we focus on visual representation learning, leaving other modalities for future exploration.
For image overlays, we sample $p$ images from a predefined dataset to form a pool for each run. The benchmark is dataset-agnostic, enabling the use of any image dataset, including images generated procedurally via diffusion. At the beginning of each episode, SPGym selects a random image from the run pool, splits it into $H\times W$ indexed patches, and overlays it onto the puzzle. The agent must then solve the puzzle by reconstructing the shuffled image, testing its ability to form compositional visual representations from raw pixel data.
Complexity Scalability
A key feature of our benchmark is its scalable difficulty. SPGym can increase representation learning complexity by keeping the grid size fixed while expanding the pool of images. This approach holds the underlying puzzle dynamics constant, ensuring that the increased difficulty comes solely from the agent’s need to handle a larger variety of visual observations.
SPGym also scales by adjusting grid sizes. Larger grids increase the search space for solving the puzzle, though the core algorithmic problem remains unchanged. [^2] The same algorithm can be used to solve the puzzle, regardless of grid size [12, 17]. Larger grid sizes also increase the complexity of representation learning by splitting images into more patches. Although scaling the grid size helps to evaluate the efficacy of policy-learning algorithms in more challenging settings, we find that a $3\times 3$ grid is sufficient to differentiate the performance of the algorithms tested in this study.
By maintaining fixed environment dynamics and task mechanics across different levels of difficulty, SPGym ensures that performance variations reflect the agent’s ability to learn and generalize visual representations, rather than adapt to changing dynamics. Consequently, SPGym rigorously assesses an agent’s capability to handle increasingly complex visual inputs while preserving consistent underlying behaviors.
@@ 4 Experimental Setup
We evaluate the performance of both model-free and model-based reinforcement learning agents using the image-based variation of SPGym. To simulate open-world conditions, we select images from the validation split of the ImageNet-1k [13] dataset. [^3] Available at https://huggingface.co/datasets/ILSVRC/imagenet-1k This choice allows future work to leverage encoders pretrained on the train split, ensuring reproducibility and comparability with our results.
For model-free experiments, we employ Proximal Policy Optimization [14, PPO] agents. In the model-based category, we use agents based on the DreamerV3 [4] architecture, which includes an autoencoder to enhance representation learning. To evaluate how well SPGym differentiates agents based on their representation learning capabilities, we compare the performance of standard PPO and DreamerV3 agents with variants that alter the representation learning process. Specifically, we include a PPO variant with pretrained encoders and a DreamerV3 variant without gradients from the decoder (by setting the decoder’s loss_scale to 00). For the pretrained PPO agents, we use encoders from agents trained on both the same pool and different pools of images. This allows us to assess, respectively, potential improvements in sample efficiency with better encoders and the generalization capabilities of these pretrained models to new visual inputs. We load encoders pretrained on the same pool sizes and leave the exploration of how encoders trained with larger pool sizes generalize to unseen pools for future work. In all configurations, gradients propagate through the entire model.
@@
Our hardware setup comprises an AMD Ryzen 7 3700X CPU, an NVIDIA RTX 3090 GPU, 64GB of RAM, and 128GB of swap space. Using this configuration, the most time-consuming DreamerV3 runs require approximately 20 hours, while the longest PPO runs complete in around 30 minutes. [^4] We attribute this runtime disparity primarily to DreamerV3’s extensive utilization of swap space for its replay buffer.
@@ 5 Results
Figure 3: Average number of steps to solve the puzzle as a function of the pool size. We consider the puzzle solved when the agents reach $80\%$ success rate and limit the maximum number of environment steps per run to 10 million. Error bars represent the $95\%$ confidence interval of a bootstrap with 1,000 resamples from 5 independent seeds.
@@
The results in Figure 3 demonstrate that as the complexity of the image pool increases, model-based agents, particularly the standard DreamerV3, consistently outperform their model-free counterparts. The DreamerV3 agent without the decoder performs noticeably worse, emphasizing the importance of DreamerV3’s autoencoder structure in representation learning. For image pools with two or more images, DreamerV3 maintains a clear advantage over PPO-based agents, which is expected given the absence of dedicated representation learning mechanisms in PPO and the numerous optimizations present in DreamerV3.
Interestingly, we observe that PPO agents with pretrained encoders perform better than those without, even when the pretraining was done on different image pools. This suggests that these encoders capture transferable, task-agnostic features that aid performance. However, as the representation learning challenge increases (i.e., larger image pools), the efficiency of all agents diminished. This trend reflects the scaling difficulty introduced by SPGym, further highlighting its utility as a benchmark for testing agents’ ability to generalize across diverse visual inputs.
Figure 4 illustrates the success rates as a function of environment steps for both standard DreamerV3 and PPO agents, across various image pool sizes. As the pool size increases, both agents exhibit a decline in sample efficiency. Even the DreamerV3 agent, which generally performs well, fails to solve the puzzle when the pool size scales, underscoring the difficulty of scaling representation learning in complex visual environments. This result demonstrates the potential of SPGym to challenge even state-of-the-art agents, revealing limitations in their ability to generalize from raw visual inputs.
Overall, our experiments show that SPGym effectively discriminates between agents based on their ability to form robust and generalizable state representations. The increasing difficulty, driven by visual diversity, highlights key challenges that remain in representation learning for decision-making tasks. These results open up several avenues for future research, particularly in developing agents that can handle more complex and diverse observation spaces.

### v2

@@ Abstract
Learning effective visual representations is crucial in open-world environments where agents encounter diverse and unstructured observations. This ability enables agents to extract meaningful information from raw sensory inputs, like pixels, which is essential for generalization across different tasks. However, evaluating representation learning separately from policy learning remains a challenge in most reinforcement learning (RL) benchmarks. To address this, we introduce the Sliding Puzzles Gym (SPGym), a benchmark that extends the classic 15-tile puzzle with variable grid sizes and observation spaces, including large real-world image datasets. SPGym allows scaling the representation learning challenge while keeping the latent environment dynamics and algorithmic problem fixed, providing a targeted assessment of agents’ ability to form compositional and generalizable state representations. Our experiments with both model-free and model-based RL algorithms, with and without explicit representation learning components, show that as the representation challenge scales, SPGym effectively distinguishes agents based on their capabilities. Moreover, SPGym reaches difficulty levels where no tested algorithm consistently excels, highlighting key challenges and opportunities for advancing representation learning for decision-making research.
@@ 1 Introduction
Learning meaningful representations from raw sensory inputs, such as visual data, is fundamental to reinforcement learning (RL) agents’ ability to generalize across different tasks in complex, open-world environments. In visual RL, agents must process high-dimensional pixel data, extract useful features, and utilize these features for decision-making. This becomes especially crucial as real-world applications demand adaptability to unstructured and diverse observations. However, measuring an agent’s representation learning capabilities independently from other learning tasks, such as policy optimization or dynamics modeling, remains a key challenge in RL benchmarks.
Widely adopted RL benchmarks like Atari [2], DeepMind Control Suite [16], and DeepMind Lab [1] focus on evaluating overall agent performance, where representation learning occurs alongside policy optimization and dynamics modeling. While specialized benchmarks have emerged to address specific aspects of visual learning — such as the Distracting Control Suite [15] for robustness to visual noise and ProcGen [3] for generalization through procedural generation — their primary goals do not include isolating representation learning capabilities. This makes it challenging to evaluate an agent’s ability to form robust state representations independently from other learning aspects.
We address this gap by introducing the Sliding Puzzles Gym (SPGym), [^1] Code available at https://github.com/bryanoliveira/sliding-puzzles-gym. a benchmark designed to isolate and evaluate visual representation learning capabilities in RL agents. Unlike existing benchmarks where representation learning intertwines with policy optimization and dynamics modeling, SPGym provides a controlled environment where the underlying task structure and dynamics remain fixed while the visual complexity can scale systematically. By extending the classic 15-tile puzzle with variable grid sizes and rich observation spaces, such as image-based tiles, SPGym enables researchers to precisely measure how well agents can form robust state representations independent of their policy learning abilities.
Our experiments evaluate the performance of both model-free and model-based RL agents on SPGym, assessing their ability to handle increasingly complex visual input spaces. We compare standard PPO [14] and DreamerV3 [4] agents with variants that modify their representation learning components. The results demonstrate that, as the image pool size increases, SPGym effectively differentiates between agents based on their representation learning capabilities. Specifically, model-based agents like DreamerV3 [4] outperform model-free ones like PPO [14] in most scenarios, though even these methods struggle with higher levels of visual complexity.
Contributions.
@@
  • We introduce the Sliding Puzzles Gym (SPGym), a scalable benchmark for assessing representation learning in visual decision-making by allowing systematic scaling of visual complexity while maintaining fixed environment dynamics;
  • Through experiments with model-free and model-based agents, we show that SPGym effectively differentiates algorithms based on their representation learning capabilities; and
  • We reveal limitations in current RL methods’ ability to handle increasingly complex and diverse visual inputs, showcasing the scalability and challenge SPGym offers as a benchmark.
Through SPGym, we provide a focused benchmark that isolates representation learning challenges from policy learning, enabling targeted research into visual RL methods. This separation allows researchers to systematically improve agents’ ability to handle complex visual inputs, a critical capability for real-world applications.
@@ 2 Related Work
Reinforcement learning benchmarks are essential tools for evaluating agent performance across various tasks. Popular benchmarks like Atari [2], DeepMind Control Suite [16], and DeepMind Lab [1] primarily assess overall agent performance, inherently combining representation learning with policy optimization and dynamics modeling. This integration makes it challenging to isolate the specific impact of representation learning on an agent’s performance. Even methods designed to enhance representation learning for RL agents, such as DARLA [5], CURL [8], RAD [9], DrQ [7], CBM [10], and CycAug [11], are typically evaluated within these entangled settings, preventing a clear, isolated assessment of their representation learning capabilities.
Among existing benchmarks, ProcGen [3] is one of the closest to our work. ProcGen takes a step towards evaluating visual generalization through diversity by offering procedurally generated levels that challenge agents to adapt to unseen environments. However, it does not provide a controlled way to isolate representation learning from policy and dynamics learning. In ProcGen, the complexity of procedurally generated levels can obscure the specific contributions of representation learning, as agents must simultaneously learn representations, policies, and environment dynamics. This entanglement makes it challenging to pinpoint whether performance improvements stem from better representations or other factors such as improved policy learning or exploration strategies.
Alternatively, the Distracting Control Suite [15], an extension of DM Control, introduces visual distractions to evaluate agents’ robustness to irrelevant variations. While it allows for parametrizable control over distraction complexity, its primary focus is on agents’ ability to ignore these distractions rather than to extract meaningful information from visually complex observations, assessing resilience to noise but not the capability to learn and utilize rich visual representations.
SPGym addresses these limitations by providing a controlled environment where difficulty arises solely from the complexity of visual observations, rather than changes in task dynamics or objectives. This design allows for a more precise evaluation of an agent’s representation learning capabilities. By systematically scaling the visual complexity while keeping the underlying task structure and dynamics fixed, SPGym enables researchers to isolate and measure the impact of representation learning independently from other learning aspects. This makes SPGym particularly well-suited for comparing RL algorithms that specifically target improvements in representation learning.
@@ 3 The Sliding Puzzles Gym
Figure 1: Overview of the Sliding Puzzles Gym (SPGym). The framework extends the 15-tile puzzle by incorporating image-based tiles, allowing scalable representation complexity while maintaining fixed environment dynamics. Agents must solve the puzzle using only image observations.
The Sliding Puzzles Gym (SPGym) extends the 15-tile puzzle, where players must rearrange a shuffled $4\times 4$ grid of numbered tiles into the correct sequence using four actions: UP, DOWN, LEFT, or RIGHT. SPGym generalizes this concept by supporting variable grid sizes, from $2\times 2$ to larger $H\times W$ configurations, and incorporating diverse observation spaces. Here, we focus on $3\times 3$ grids to emphasize visual representation learning, providing only image-based observations to the agent.
Environment Dynamics and Actions.
At each step, the agent can slide a tile adjacent to the empty space into that position, making the action space discrete and deterministic, as illustrated in Figure 1. The objective is to rearrange shuffled image tiles into their correct positions, forming a complete, intelligible image. While the underlying puzzle mechanics remain simple and consistent, the visual representation challenge scales with the diversity of images used. Agents must learn to extract meaningful features from potentially complex visual inputs, recognize patterns across different images, and understand how these features relate to the puzzle’s solution state. This separation between fixed dynamics and scalable visual complexity allows SPGym to isolate and evaluate an agent’s representation learning capabilities independently from its policy learning abilities.
Observation Spaces and Wrappers.
SPGym’s internal representation is a 2D array of tile indices representing the current grid state. We provide flexible wrappers that convert states into various observation modalities, ranging from simple one-hot encodings to complex visual observations, such as real-world images and text (Figure 2). This adaptability makes SPGym a versatile benchmark for exploring how reinforcement learning agents handle diverse observation spaces. In this paper, we focus on visual representation learning, leaving other modalities for future exploration.
For visual observations, we employ image overlays. In each run, we sample $p$ images from a predefined dataset to form a pool. SPGym is dataset-agnostic, allowing the use of any image dataset, including procedurally generated images. At the start of each episode, we select a random image from the pool, split it into $H\times W$ indexed patches, and overlay it onto the puzzle. The agent’s task is to reconstruct the shuffled image, testing its ability to form compositional representations from pixels.
To initialize the puzzle’s state, SPGym offers two methods. The primary method, used in this paper, generates a random $H\times W$ array and ensures solvability by adjusting the parity of the puzzle, if necessary, by swapping the first two tiles [6]. The second method, which can facilitate curriculum learning in future work, begins with a solved puzzle and applies a series of random valid moves to create an initial state, allowing for the selection of easier starting configurations.
Reward Function.
At each time step, the reward is computed as follows:
$r_{t}=\begin{cases}-\frac{\sum_{i,j}|x_{i,j}-x^{*}_{i,j}|+|y_{i,j}-y^{*}_{i,j}|}{D},&\text{if action is valid}\\ {\color[rgb]{0.5,0.0,0.0}-1,}&\text{if action is invalid}\\ {\color[rgb]{0.0,0.3,0.0}+10,}&\text{if puzzle is solved}\end{cases}\text{, with}$ (1)
@@
Complexity Scalability.
A key feature of our benchmark is its scalable difficulty. SPGym increases representation learning complexity by keeping the grid size fixed while expanding the pool of images. This approach holds the underlying puzzle dynamics constant, ensuring that the increased difficulty comes solely from the agent’s need to handle a larger variety of visual observations.
SPGym also scales the challenge by adjusting grid sizes. Larger grids increase the search space for solving the puzzle, though the core algorithmic problem remains unchanged. [^2] The same algorithm can solve the puzzle regardless of grid size [12, 17]. Larger grid sizes also increase the complexity of representation learning by splitting images into more patches. Although scaling the grid size helps evaluate the efficacy of policy-learning algorithms in more challenging settings, we find that a $3\times 3$ grid sufficiently differentiates the performance of tested algorithms.
By maintaining fixed environment dynamics across different difficulty levels, SPGym ensures that performance variations reflect the agent’s ability to learn and generalize visual representations, rather than adapt to changing dynamics. Consequently, SPGym rigorously assesses an agent’s capability to handle increasingly complex visual inputs while preserving consistent underlying behaviors.
@@ 4 Experimental Setup
We evaluate the performance of both model-free and model-based reinforcement learning agents using the image-based variation of SPGym. To simulate open-world conditions, we select images from the validation split of the ImageNet-1k [13] dataset. [^3] Available at https://huggingface.co/datasets/ILSVRC/imagenet-1k. This choice allows future work to leverage encoders pretrained on the train split, ensuring reproducibility and comparability with our results.
Agents and Variants.
For model-free experiments, we employ Proximal Policy Optimization [14, PPO] agents. In the model-based category, we use agents based on the DreamerV3 [4] architecture, which includes an autoencoder to enhance representation learning. To evaluate how well SPGym differentiates agents based on their representation learning capabilities, we compare the performance of standard PPO and DreamerV3 agents with variants that alter the representation learning process. Specifically, we include a PPO variant with pretrained encoders and a DreamerV3 variant without gradients from the decoder (by setting the decoder’s loss_scale to 00). For the pretrained PPO agents, we use encoders from agents trained on both the same pool and different pools of images. This allows us to assess potential improvements in sample efficiency with better encoders and the generalization capabilities of these pretrained models to new visual inputs. We load encoders pretrained on the same pool sizes and leave the exploration of how encoders trained with larger pool sizes generalize to unseen pools for future work. In all configurations, gradients propagate through the entire model.
Experimental Details.
@@
Our hardware setup comprises an AMD Ryzen 7 3700X CPU, an NVIDIA RTX 3090 GPU, 64GB of RAM, and 128GB of swap space. Using this configuration, the most time-consuming DreamerV3 runs require approximately 20 hours, while the longest PPO runs complete in around 30 minutes. [^4] This runtime disparity is primarily due to DreamerV3’s extensive use of swap space for its replay buffer.
Figure 3: Number of environment samples to solve the puzzle as a function of the pool size (lower is better). We consider the puzzle solved when the agents reach $80\%$ success rate and cap the number of environment steps per run to 10 million. Error bars represent the $95\%$ confidence interval, calculated using bootstrap resampling with 1,000 iterations from 5 independent seeds.
@@ 5 Results
In this section, we address the following key research questions through our experiments:
Can SPGym distinguish agents based on their representation learning performance?
Figure 3 demonstrates that SPGym effectively differentiates agents based on their representation learning capabilities. Model-based agents, especially the standard DreamerV3, consistently outperform model-free agents as the image pool complexity increases. This indicates that SPGym not only assesses the agents’ task-solving abilities but also their skill in learning and applying robust state representations. The significant performance gap underscores the critical role of advanced representation learning mechanisms in handling complex environments.
How do representation-learning specific components impact agent performance in SPGym?
Our analysis indicates that representation-learning components benefit agent performance. The DreamerV3 agent, with its integrated autoencoder structure, outperforms its variant without the decoder. This underscores the role of autoencoders in enhancing the agent’s ability to abstract and generalize from visual inputs, which is crucial in environments with high visual complexity. The presence of these components allows agents to form more nuanced and effective representations, directly impacting their decision-making efficiency.
Does pretraining encoders improve the performance of model-free agents?
Pretraining encoders significantly enhances the performance of PPO agents, enabling them to outperform their non-pretrained counterparts, even when the pretraining occurs on different image pools. This suggests that pretrained encoders capture essential, transferable features that benefit learning across various tasks. It highlights the potential of leveraging pretraining as a strategy to enhance the adaptability and efficiency of model-free agents in diverse environments.
How does the complexity of the image pool affect the performance of different agents?
Figure 4 shows that increasing the image pool size leads to a decline in sample efficiency for both DreamerV3 and PPO agents. This trend highlights the challenges in scaling representation learning for more complex visual inputs. Even DreamerV3, which generally excels, struggles with larger pool sizes, indicating that current techniques may not fully capture the intricacies of diverse environments. This underscores SPGym’s potential to push the boundaries of current methodologies and inspire more advanced representation learning strategies.
Our experiments confirm that SPGym effectively differentiates agents based on their representation learning capabilities. The benchmark’s increasing difficulty, driven by visual diversity, highlights ongoing challenges in representation learning for decision-making tasks. These results suggest promising directions for future research, particularly in developing agents capable of handling more complex and diverse observation spaces.